OpenStudy (anonymous):

Brian has a bag that contains 14 red marbles and 12 yellow marbles. He selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Brian selects a red marble and then a yellow marble? Round your answer to the nearest percent. P(red and yellow)

OpenStudy (anonymous):

26%

OpenStudy (anonymous):

46%

OpenStudy (anonymous):

The correct answer is 26% because 14/26 * 12/25 = 0.25846154.

OpenStudy (anonymous):

26%

OpenStudy (michele_laino):

probability to get a red marble is 14/26 whereas the probability to get an yellow marble is 12/25 I supposed to get a red marble for first. So the requested probability is: \[\huge p = \frac{{14}}{{26}} \times \frac{{12}}{{25}} \times 100 = ...?\]

OpenStudy (anonymous):

so 26%

OpenStudy (michele_laino):

yes! We get \(25.8 \%\) which can be rounded to \(26\%\)

OpenStudy (anonymous):

k

OpenStudy (michele_laino):

:)